Get Help

Online: Request Help
Phone: (707) 826-4357 • Hours
Walk-In: Library 101 • Hours
Reset HSU Password
System Status

System Status Archive

System Issue start Who is affected What happened
Other Aug 14 2016 - 05:00 All users Maintenance Window has been complete. Nolij is still migrating but will return today. Aug 12 2016 - 11:01 Anyone logging into a campus computer who enters their password incorrectly 3 times within a minute are locked out of their account for 5 minutes. To reduce the inconvenience for everyone, while still maintaining security, we will be changing this to 6 incorrect entries in a minute will result in a 5 minute lock out. **Update: we have changed the number of unsuccessful login attempts in 1 minute resulting in a 5 min logout to 6.
Off-Campus Internet Aug 11 2016 - 05:00 People or devices using the campus internet connection Regular maintenance on Border Firewall. **UPDATE** Maintenance complete, ~15 seconds of downtime was incurred.
PeopleSoft HCM Aug 10 2016 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, or Admissions. During our regularly scheduled maintenance window the following items will be installed in PeopleSoft production, Advising Center Fields Additions (HUM_FC419 modification), MP 21.02 and Oracle POC 928675 (technical fix to address database latching/CPU issues during high use of Remote COBOLs).
Other Aug 4 2016 - 05:45 All users of shibboleth SSO services including OBI, CSU portal, zimride, zoom. Restarting service. Completed restart.
Other Aug 1 2016 - 13:11 All Campus Systems That Use Storage Planned maintenance and upgrading the Principal switches on the Storage Area Network. *UPDATE 2 August 1045- fabric upgrade is progressing and no problems have been encountered. Any issues with server/service availability during the process should be reported to the Help Desk at x4357 or
Google Apps - Mail Jul 28 2016 - 08:00 People connecting to email using the imap protocol IMAP was being disabled in people's accounts via our new account management tool. We have changed the setting and are re-enabling the setting. We should have it resolved by the end of the day.
Account Settings Aug 1 2016 - 01:00 About 300 people who's passwords expired early Monday morning. There is a bug in our new account management system that we are investigating. We are hoping to have a resolution for this in place by the end of Tuesday.
Campus Data Network Aug 1 2016 - 16:45 Library data network During a switch replacement in Science A, cisco UDLD datagrams were received at the Library switch. This caused the Library switch to disable its uplink for a few minutes.
Campus Data Network Aug 1 2016 - 16:10 It is possible that Network Users at Science A will lose network connectivity - both wired and wireless. TNS is replacing some failing equipment before anything can go wrong. We do not expect an outage, but if one does occur, it will be approximately 15 to 20 minutes. *UPDATE - Faulty network switch replaced.